# SelfSplat: Pose-Free and 3D Prior-Free Generalizable 3D Gaussian Splatting

### v3

@@ Abstract
We propose SelfSplat, a novel 3D Gaussian Splatting model designed to perform pose-free and 3D prior-free generalizable 3D reconstruction from unposed multi-view images. These settings are inherently ill-posed due to the lack of ground-truth data, learned geometric information, and the need to achieve accurate 3D reconstruction without finetuning, making it difficult for conventional methods to achieve high-quality results. Our model addresses these challenges by effectively integrating explicit 3D representations with self-supervised depth and pose estimation techniques, resulting in reciprocal improvements in both pose accuracy and 3D reconstruction quality. Furthermore, we incorporate a matching-aware pose estimation network and a depth refinement module to enhance geometry consistency across views, ensuring more accurate and stable 3D reconstructions. To present the performance of our method, we evaluated it on large-scale real-world datasets, including RealEstate10K, ACID, and DL3DV. SelfSplat achieves superior results over previous state-of-the-art methods in both appearance and geometry quality, also demonstrates strong cross-dataset generalization capabilities. Extensive ablation studies and analysis also validate the effectiveness of our proposed methods. Code and pretrained models are available at https://gynjn.github.io/selfsplat/
@@ 2.1 Pose-free Neural 3D Representations
In the absence of camera pose information, recent efforts have aimed to jointly optimize camera poses and 3D scenes. Starting with optimization-based methods, BARF [30] and subsequent research [2, 16, 24] addressed this challenge by training poses along with implicit or explicit scene representations. Also, in a generalizable setting with NeRF representations, FlowCam [42] utilized pretrained flow estimation model, RAFT [46], and find the rigid-body motion between 3D point clouds using the Procrustes algorithm [11]. DBARF [6] extended the previous optimization-based method [30] and utilized recurrent GRU [10] network for pose and depth estimation. Based on 3D-GS, GGRT [29] showed first pose-free generalizable method by employing self-supervised learning. However, these methods face practical limitations due to their reliance on pretrained models [42], the need for additional fine-tuning stages [6, 29], and the computationally intensive volume rendering process [6, 42], all of which hinder their scalability and efficiency in real-world applications. Also, CoPoNeRF [21] provides poses and radiance fields estimation at the inference stage, it still requires ground-truth pose supervision during training. In contrast, our method can reconstruct 3D scenes and synthesize novel views from unposed images, mitigating the preceding challenges, and offering a more scalable and efficient solution.
@@ 5.1 Experiment Setup
Evaluation metrics. For novel view synthesis, we use standard evaluation metrics: PSNR, SSIM, and LPIPS. Pose estimation is evaluated based on geodesic rotation and translation angular error, following the approach in [21]. For RE10k and ACID, we categorize test set context pairs by image overlap ratios to evaluate performance across small (0.05-0.6%), medium (0.6-0.8%), and large (0.8%+) overlap, identified by a pretrained dense image matching method [14]. For DL3DV, overlap categories are defined by frame intervals between context images: 6 frames for large and 10 frames for small overlap.
Implementation details. We employ the encoder part of pretrained CroCo [54] model as our monocular encoder, which is trained in a self-supervised manner, and utilized adapter [5] block designed to efficiently adapt pretrained ViT models to downstream tasks. For the Gaussian rasterizer, we implement it using gsplat [60], an open-source library for Gaussian Splatting [25], offering efficient computation and memory usage. We train RE10k and ACID with 256 $\times$ 256 resolution, and for DL3DV, we use 256 $\times$ 448 to accommodate the wider view in our experiments. For additional details, consult the supplementary material.
@@ 6 Conclusion
Limitations. While we demonstrate high-quality 3D geometry estimation in this work, the current framework still possesses limitations. First, further technical improvements are needed to support wider baseline scenarios, such as a 360 scene reconstruction from unposed images in a single forward pass. Second, our framework struggles with dynamic scenes where both camera and object motion are present. Addressing these complex scenarios may benefit from incorporating multi-modal priors [40, 44] for robust and consistent alignment across wide and dynamic scene space.

### v4

@@ Abstract
We propose SelfSplat, a novel 3D Gaussian Splatting model designed to perform pose-free and 3D prior-free generalizable 3D reconstruction from unposed multi-view images. These settings are inherently ill-posed due to the lack of ground-truth data, learned geometric information, and the need to achieve accurate 3D reconstruction without finetuning, making it difficult for conventional methods to achieve high-quality results. Our model addresses these challenges by effectively integrating explicit 3D representations with self-supervised depth and pose estimation techniques, resulting in reciprocal improvements in both pose accuracy and 3D reconstruction quality. Furthermore, we incorporate a matching-aware pose estimation network and a depth refinement module to enhance geometry consistency across views, ensuring more accurate and stable 3D reconstructions. To present the performance of our method, we evaluated it on large-scale real-world datasets, including RealEstate10K, ACID, and DL3DV. SelfSplat achieves superior results over previous state-of-the-art methods in both appearance and geometry quality, also demonstrates strong cross-dataset generalization capabilities. Extensive ablation studies and analysis also validate the effectiveness of our proposed methods.
@@ 2.1 Pose-free Neural 3D Representations
In the absence of camera pose information, recent efforts have aimed to jointly optimize camera poses and 3D scenes. Starting with optimization-based methods, BARF [33] and subsequent research [2, 17, 26] addressed this challenge by training poses along with implicit or explicit scene representations. Also, in a generalizable setting with NeRF representations, FlowCam [46] utilized pretrained flow estimation model, RAFT [51], and find the rigid-body motion between 3D point clouds using the Procrustes algorithm [11]. DBARF [6] extended the previous optimization-based method [33] and utilized recurrent GRU [10] network for pose and depth estimation. Based on 3D-GS, several methods [32, 23, 66] showed pose-free generalizable method by employing explicit 3D representation. However, these methods face practical limitations due to their reliance on pretrained models [6, 46, 32, 66], the need for additional fine-tuning stages [6, 32], and the computationally intensive volume rendering process [6, 46], all of which hinder their scalability and efficiency in real-world applications. Also, CoPoNeRF [22] provides poses and radiance fields estimation at the inference stage, it still requires ground-truth pose supervision during training. In contrast, our method can reconstruct 3D scenes and synthesize novel views from unposed images, mitigating the preceding challenges, and offering a more scalable and efficient solution.
@@ 5.1 Experiment Setup
Evaluation metrics. For novel view synthesis, we use standard metrics: PSNR, SSIM [57], and LPIPS [71]. Pose estimation is evaluated based on geodesic rotation and translation angular error, following [22]. For RE10k and ACID, we categorize test context pairs by image overlap ratios to evaluate performance across small(0.05-0.6), medium(0.6-0.8), and large(0.8+) overlap, identified by a pretrained image matching method [15]. For DL3DV, overlap categories are defined by frame intervals between context images: 6 frames for large and 10 frames for small overlap.
Implementation details. We employ the encoder part of pretrained CroCo [59] model as our monocular encoder, which is trained in a self-supervised manner, and utilized adapter [5] block designed to efficiently adapt pretrained ViT models to downstream tasks. For the Gaussian rasterizer, we implement it using gsplat [67], an open-source library for Gaussian Splatting [27], offering efficient computation and memory usage. We train RE10k and ACID with 256 $\times$ 256 resolution, and for DL3DV, we use 256 $\times$ 448 to accommodate the wider view in our experiments.
@@ 6 Conclusion
[…]
Supplementary Material
Appendix A Additional Details
A.1 Architectural Details
For the prediction of 3D Gaussians [27], we utilize the monocular, multi-view encoder and the fusion block. Unlike previous methods that utilize DepthAnything [64] as a monocular encoder [70, 62] or UniMatch [61] as a multi-view encoder [7, 47], we only employ the encoder part of Croco [58] as our monocular encoder which is trained in a fully self-supervised manner. For the multi-view encoder, we adopt the backbone of [61] with randomly initialized weights. Then, we unify features from monocular and multi-view encoders using DPT [40] block. For a detailed architecture for the fusion module, see Fig. 7.
Figure 7: Detailed 3D Gaussian prediction architecture. This module takes only context images as input.
A.2 Implementation Details
For our monocular encoder, we utilized Adapter [5], which keeps the model parameters frozen while training additional residual networks for each layer. Specifically, a residual MLP block, comprising a down-projection layer and an up-projection layer, is introduced within each layer of the transformer encoder. Considering the channel dimension of the original encoder, $C^{\text{mono}}=1024$, we set the low rank hidden dimension of AdaptMLP, $C^{\text{adapt}}=32$, to efficiently reduce computational overhead while maintaining sufficient capacity for adaptation.
For 3D Gaussian primitives, we set the order of spherical harmonics expansion to 1, enabling the representation to extend beyond the Lambertian color model. When warping the color model from each frame’s local coordinate system into an integrated global space which requires the Wigner matrices in general case, we simplify the rotation of the first level of spherical harmonics, $Y_{1}(r_{d})=[Y_{1}^{-1}(r_{d}),Y_{1}^{0}(r_{d}),Y_{1}^{1}(r_{d})]$, as follows:
$Y_{1}(r_{d})=\sqrt{\frac{3}{4\pi}}\Pi\hskip 1.42262ptr_{d},\quad\Pi=\begin{bmatrix}0&1&0\\ 0&0&1\\ 1&0&0\\ \end{bmatrix},$
where $r_{d}\in\mathbb{S}^{2}$ is the viewing direction derived from the estimated camera poses. We adopt this warping protocol from Splatter Image [50] which is a pose-required generalizable 3D reconstruction model using 3D Gaussian Splatting.
A.3 Training Details
We train all baseline models, including ours, using custom data loaders. For RealEstate10K [75] (RE10k) and ACID [36] datasets, the distance between context frames is progressively increased from 5 to 25, and target frames are randomly selected between the context frames within this range. Each model is trained for 200K iterations and for baselines we used the default hyperparameter settings provided by the respective authors. The only exception is DBARF [6], which is trained for 400K iterations due to its official implementation supporting only a batch size of one. We provide our detailed training hyperparameters in Tab. 9 and we train our model on a singe H100 GPU, which takes approximately for 3 days. For the experiment on DL3DV [35] dataset, we initialize the model with pretrained weights from RE10k dataset and train it for 50K iterations on a single H100 GPU with a batch size of 6. The distance between context frames is gradually increased from 2 to 10. This procedure is applied to FlowCAM [46] in the same way which is the baseline model on DL3DV dataset.
For VAE [29], which was initially designed for novel view synthesis from a single image, we modify its architecture following the approach in [69] to handle multi-view input images. Specifically, we employ two separate encoders and use their mean output as the input to the decoder which synthesize novel view images. All other hyperparameters remain the same as the official implementation.
SelfSplat
Config Value
optimizer Adam [28]
scheduler Linear
learning rate 1e-4
gradient clipping 0.5
batch size 12
total iters. 200,000
warmup iters. 2,000
Table 9: Training hyperparameters.
A.4 Evaluation Details
During the evaluation on RE10k and ACID datasets, we set the interval between context frames to 40 and select the middle frame as the target view point. This target frame is used as the ground truth for metric evaluations in novel view synthesis and camera pose estimation. For the overlap categories, we utilize the pretrained feature matching model, RoMa [15], to estimate the overlap ratios between the first context frame and the target frame.
For RE10k dataset, the split proportions are 18.26% for large, 60.56% for medium, and 21.17% for small categories. In ACID dataset [36], the proportions are 33.05% for large, 41.15% for medium, and 25.80% for small.
Appendix B Additional Experiment Analysis
B.1 Inference Cost
We report the memory and time consumption required to synthesize a single 256 × 256 image during the inference stage in Table 10. Memory usage is measured as the peak memory during inference, while the number of rays per batch is adjusted if necessary. Except for VAE [29], which generates novel view images without rendering operations (utilize 2D CNN blocks) and thus fail to reconstruct interpretable 3D scene representations, our method achieves significantly lower memory usage and faster rendering speed with explicit 3D representations, demonstrating its efficiency and practical usage in real-world scenarios.
Method Mem. (GB) Time (s)
VAE [29] 4.694 0.0003
DBARF [6] 2.079 0.254
FlowCAM [46] 16.644 0.801
CoPoNeRF [22] 16.802 5.624
Ours 1.795 0.002
Table 10: Memory and time consumption analysis. All baselines including ours are measured on a single NVIDIA RTX 4090 GPU.
B.2 Using N Context Views
We further evaluate the model’s performance across various numbers of input views, considering its practical application where more than two views are commonly used. The total number of frames is evenly divided based on the number of context views, and target frames are sampled between the context frames. Additionally, we generate a camera trajectory using the selected view points (context and target), and the Absolute Trajectory Error (ATE) is measured to validate the accuracy of the reconstructed camera path. We evaluate on RE10k dataset with 3 context views (80 frames) and 4 context views (120 frames) settings. As shown in Tab. 11 and Fig. 8, our method demonstrates superior performance in both novel view synthesis and camera trajectory estimation, as well as its ability to scale effectively with multiple input views and estimations over extended ranges without any further finetuning.
3 views (80 frames) 4 views (120 frames)
[…]
FlowCAM [46] 19.75 0.630 0.412 0.048 18.91 0.606 0.449 0.081
Ours 21.12 0.761 0.241 0.031 19.52 0.717 0.283 0.053
Table 11: Quantitative results of using different numbers of context views on RE10k dataset.
Figure 8: Visualization of camera trajectory on RE10k dataset. Construction of trajectory only consider the translation part of the estimated camera poses.
B.3 Additional Comparison
For the reader’s reference, we provide a comparison with Splatt3R [45], which is also a pose-free, feed-forward Gaussian Splatting method for 3D reconstruction and novel view synthesis from stereo pairs. We omitted this model in the main paper because it requires ground-truth depth and camera pose annotations during training, which are not available in the datasets we used: RE10k, ACID [36], and DL3DV [35]. Acknowledging the differences in training data—Splatt3R was trained on ScanNet++ [12], whereas our model was trained on RE10k—we evaluate them on the DTU [24] dataset, which is an out-of-distribution dataset for both. As shown in Tab. 12 and Fig. 9, our method achieves better performance than the baseline in both evaluation metrics and visual quality, and also outperforms pixelSplat [3] which is a pose-required method in training and evaluation stage. The main reason Splatt3R cannot estimate a consistent scene scale is its reliance on a fixed pretrained MASt3R [31] model, which is trained using metric camera poses, and difference between estimated intrinsic parameters and ground truth intrinsic parameters. Thus, using the DTU dataset, which consists of unseen novel scenes, they fail to align consistent 3D Gaussians.
[…]
ScanNet++ [12] Splatt3R [45] 10.24 0.295 0.629
pixelSplat [3] 12.89 0.382 0.560
RE10k [75] MVSplat [7] 13.94 0.473 0.385
Ours 13.14 0.425 0.448
Table 12: Quantitative results of novel view synthesis on DTU dataset. While pixelSplat [3] and MVSplat [7] are pose-required methods, we include them for the reader’s reference.
Figure 9: Qualitative comparison of novel view synthesis on DTU dataset.
B.4 Baseline Comparisons
We provide additional baseline results on cross-dataset generalization in Tab. 13.
RE10k $\rightarrow$ ACID ACID $\rightarrow$ RE10k
Method PSNR $\uparrow$ SSIM $\uparrow$ ${}^{\circ}\downarrow$ ${}^{\circ}\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ ${}^{\circ}\downarrow$ ${}^{\circ}\downarrow$
VAE 23.67 0.649 1.619 117.721 18.98 0.537 3.744 65.884
DBARF 23.56 0.644 1.772 51.969 12.60 0.502 3.306 47.851
Ours 26.60 0.793 1.119 18.607 21.65 0.728 1.618 17.993
Table 13: Additional comparison on cross-dataset generalization.
B.5 Additional Ablation and Analysis
We provide additional ablation studies and analyses, focusing on our encoder module. All methods are trained on RE10k [75] for 50k iterations, following the same procedure as in the main paper. As shown in Tab. 14, our feature fusion module with CroCo [58] initialization shows superior results in evaluation metrics.
Ours 22.65 0.764 0.222 1.036 13.705
No CroCo [58] Init 22.15 0.738 0.240 1.091 14.042
No Monocular Encoder 21.88 0.733 0.247 1.394 17.125
No Multi-view Encoder 21.47 0.731 0.243 1.233 16.327
Table 14: Ablation studies on the encoder module design.
Pretrained weight. Since our goal is to use only unposed raw video datasets without 3D priors, we utilized CroCo, trained in a self-supervised manner. While DUSt3R [56] or MASt3R [31] pre-trained weight could enhance performance, we focus on demonstrating that 3D foundation models can be trained without costly 3D annotations.
B.6 Architectural and Evaluation Design
We designed our evaluation protocol assuming that there are no given poses, so we made a separate pose block (context and target) and a Gaussian branch (only context) independently. Thus, target images are used to estimate camera poses for following novel view synthesis evaluations. All baselines follow this protocol in their original implementations, except for CoPoNeRF [22] which utilizes given camera poses, so we substitute these poses in CoPoNeRF with estimated ones for a fair comparison.
B.7 Depth Visualization
We also provide the visualization of depth maps generated through rendering, which is essential for producing interpretable 3D representations. By comparing the results of our method with previous approaches, as shown in Fig. 10, SelfSplat demonstrates robust and reliable depth maps derived from 3D scene structures.
Figure 10: Qualitative comparison of depth visualization on RE10k dataset. Depth maps are obtained following the rendering process.
Appendix C Limitations
@@
While we demonstrate high-quality 3D geometry estimation in this work, the current framework still possesses limitations. First, further technical improvements are needed to support wider baseline scenarios, such as a 360 scene reconstruction from unposed images in a single forward pass. Second, our framework struggles with dynamic scenes where both camera and object motion are present. Addressing these complex scenarios may benefit from incorporating multi-modal priors [43, 49] for robust and consistent alignment across wide and dynamic scene space.
Appendix D Additional Results
We provide additional results on the following pages including novel view synthesis and epipolar line visualizations.
Figure 11: Qualitative comparison of novel view synthesis on RE10k dataset.
Figure 12: Qualitative comparison of novel view synthesis on ACID dataset.
Figure 14: Epipolar lines visualization on RE10k dataset. We draw the lines from reference to target frame using relative camera pose.